i took a 45 question test and missed 5 what is my score

Answer by jim_thompson5910(35256)

You got 45-5 = 40 questions correct out of 45

So your score as a percentage is


[(number correct)/(number total)] * 100 = (40/45)*100 = 0.889*100 = 88.9%


So you got an 88.9%